# “It’s Always a Losing Game”: How Workers Understand and Resist Surveillance Technologies on the Job

### v1

@@ 5.2. The Promise and Limitations of Policy
When it came to regulation, during our interviews, workers expressed mixed feelings. While there were some workers like P1 who believed that governmental regulation could offer workers stronger data and privacy protections, as well as prevent more intrusive forms of monitoring, others were more skeptical. P6’s belief that in the United States “companies own the politicians” speaks to the incredible lobbying power corporations have at all levels of government. In 2023, the tech companies Google, Meta, Amazon, and Microsoft collectively spent 64 million dollars on lobbying at the federal level (Henshall, 2024). At the state level, lobbying groups have had multiple successes in their fight against new privacy bills. When Oregon lawmakers tried to pass a significant digital privacy law that would allow citizens to sue companies in violation in 2023, the lobbying group TechNet successfully killed the bill. Further, Calacci and Stein (Calacci and Stein, 2023) highlight how current data protection laws fail to protect workers because they focus on protecting the individual data subject, rather than protecting collective rights. While there has been growing interest at the US federal level to introduce stronger worker protections, there are currently no federal laws in place (anon, 2023; Thornburg, 2024; SCIENCE and POLICY, 2023). We must acknowledge that most meaningful policy change for workers remains a slow and uphill battle.
The experiences of international workers such as P5 and P8 shed further light on the limitations of regulation as more jobs are ”offshored” to countries with cheaper labor and weaker labor protections (Bureau, 2008). Since P5 worked remotely for a US-based company, she believed that she did not have the same labor protections as those who worked for Mexico-based companies because her employer did not have to comply with Mexican labor laws. P8 felt little hope for stronger privacy regulation in Philippines because, in her eyes, the country lacked a “solid foundation regarding security” in the first place. As work becomes more fissured and more US-based companies continue to use outsourced contract labor from countries like India and the Philippines for their customer support and data work (Tan and Cabato, 2023), this brings new questions about the role of HCI in policy to protect workers across the global supply chain. While we see great promise in the role of worker-centered design to inform governmental policy at the local, state, and even federal level (Yang et al., 2024), it is certainly not a panacea.
@@ 5.3.1. Countering Surveillance through Refusal
Through these everyday acts of resistance, workers creatively and ingenuously repurposed tools, subverted expectations, and developed new methods to reclaim their autonomy. Drawing from the radical traditions of Luddism and unmaking, we echo Sabie et al.(Sabie et al., 2023)’s call for HCI researchers to take an emancipatory approach to worker-centered design — one that draws inspiration from the existing approaches workers are already using to dismantle, disarm, and destabilize systems of control. This approach could involve designing tools and mechanisms that enable workers to obfuscate, disrupt, or neutralize WSTs, as well as helping workers to safely enact ”soldiering” practices that evade detection. More crucially, taking an emancipatory approach would go beyond merely redesigning WSTs to mitigate harms and toward the possibility of their abandonment when they fail to support workers (Johnson et al., 2024).
@@ 5.3.2. Building toward Collective Action
Rather than largely fixing our gaze, as a scholarly community, on the potential for redesigning surveillance platforms or shifting managerial practice, we can support these more extended forms of worker community building that scaffold political consciousness, circulate tactics of subversion, and prefigure more sustainable and trusting working relations. Unlike technical interventions that might seek to mitigate the most outsized harms caused by intrusive surveillance, this approach would mean the slow and intentional work of collective transformation and cultural change. It is only through this recommitment of meeting workers where they are and sharing in mutual learning, that we can finally unmake the surveillance apparatus.
@@ 6. Conclusion
In this paper, we have sought to understand the effects of the rapid expansion of workplace surveillance technologies that followed the mass move to remote work during the COVID-19 pandemic. In our analysis of anonymous posts on Reddit and of ten interviews with workers from a range of industries, we found that constant and layered surveillance created major privacy concerns for workers, destroyed trust between supervisors and supervisees, and led to significant anxiety and, ultimately, burnout. By understanding how workers responded to and resisted these effects, we call on CSCW researchers to adopt more emancipatory approaches to worker-centered-design that center on resistance and refusal.

### v2

@@ 5.2. The Promise and Limitations of Policy
When it came to regulation, during our interviews, workers expressed mixed feelings. While there were some workers like P1 who believed that governmental regulation could offer workers stronger data and privacy protections, as well as prevent more intrusive forms of monitoring, others were more skeptical. Although P2, a disabled worker, received accommodations that allowed them to take more breaks under surveillance, they felt that disability laws ultimately failed to protect them from being demoted and having their contract terminated as a consequence of requesting these accommodations. Additionally, P6’s belief that “companies own the politicians” in the United States speaks to the incredible lobbying power corporations have at all levels of government. For example, at the time of writing, the Consumer Financial Protection Bureau has introduced new guidance designed to protect workers from digital surveillance (CFPB, 2024), but is also facing increased scrutiny and calls for abolishment from powerful industry actors (Stempel, 2025; Stratford, 2024).
State privacy laws such as the California Consumer Privacy Act (CCPA) introduce important protections granting workers the right to know when employers are monitoring them and for what purpose, as well as the ability to access to their personal data, request its deletion, and limit its sale (Feng, 2023). However, Calacci and Stein (Calacci and Stein, 2023) highlight how current data protection laws such as CCPA fail to protect workers as a whole because they focus on protecting the individual data subject, rather than protecting collective rights. While there has been growing interest at the US federal level to introduce stronger worker protections, there are currently no federal laws in place (of Former Sen. Bob Casey, 2023; Congress, 2024; SCIENCE and POLICY, 2023). We must acknowledge that most meaningful policy change for workers remains a slow and uphill battle.
The experiences of international workers such as P5 and P8 shed further light on the limitations of regulation as more jobs are ”offshored” to countries with cheaper labor and weaker labor protections (Bureau, 2008). Since P5 worked remotely for a US-based company, she believed that she did not have the same labor protections as those who worked for Mexico-based companies because her employer did not have to comply with Mexican labor laws. P8 felt little hope for stronger privacy regulation in the Philippines because, in her eyes, the country lacked a “solid foundation regarding security” in the first place. As work becomes more fissured and more US-based companies continue to use outsourced contract labor from countries like India and the Philippines for their customer support and data work (Tan and Cabato, 2023), this brings new questions about the role of HCI in policy to protect workers across the global supply chain. While we see great promise in the role of worker-centered design to inform governmental policy at the local, state, and even federal level (Yang et al., 2024), it is certainly not a panacea.
@@ 5.3.1. Countering Surveillance through Refusal
Through these everyday acts of resistance, workers creatively and ingeniously repurposed tools, subverted expectations, and developed new methods to reclaim their autonomy. Drawing from the radical traditions of Luddism and unmaking, we echo Sabie et al.’s (Sabie et al., 2023) call for HCI researchers to take an emancipatory approach to worker-centered design — one that draws inspiration from the existing approaches workers are already using to dismantle, disarm, and destabilize systems of control. This approach could involve designing tools and mechanisms that enable workers to obfuscate, disrupt, or neutralize WSTs, as well as helping workers to safely enact ”soldiering” practices that evade detection. More crucially, taking an emancipatory approach would go beyond merely redesigning WSTs to mitigate harms and toward the possibility of their abandonment when they fail to support workers (Johnson et al., 2024).
@@ 5.3.2. Building toward Collective Action
Rather than largely fixing our gaze, as a scholarly community, on the potential for redesigning surveillance platforms or shifting managerial practice, we can support these more extended forms of worker community building that scaffold political consciousness, circulate tactics of subversion, and prefigure more sustainable and trusting working relations. Unlike technical interventions that might seek to mitigate the most outsized harms caused by intrusive surveillance, this approach would mean the slow and intentional work of collective transformation and cultural change. It is only through this recommitment to meeting workers where they are and sharing in mutual learning that we can finally unmake the surveillance apparatus.
@@ 6. Conclusion
In this paper, we have sought to understand the effects of the rapid expansion of workplace surveillance technologies that followed the mass move to remote work during the COVID-19 pandemic. In our analysis of anonymous posts on Reddit and of ten interviews with workers from a range of industries, we found that constant and layered surveillance created major privacy concerns for workers, deteriorated trust between supervisors and supervisees, and led to significant anxiety and, ultimately, burnout. In understanding how workers responded to and resisted these effects, we call on CSCW researchers to adopt more emancipatory approaches to worker-centered-design that center on resistance and refusal.
7. Acknowledgements
We would like to thank our interview participants who generously contributed their insights, as well as the Reddit moderators who expressed support for this project. We also thank our reviewers and colleagues who provided valuable feedback, including Hunter Akridge, Wilneida Negrón, Tucker Rae-Grant, Matthew Scherer, and Franchesca Spektor.